alis:

A and B are two events. Given that P(A)=0.75 , P(B)=0.2 and P(AandB)=0.15 .

alis:

alis:

@Vocaloid

sillybilly123:

if \(P ( A \text{ and } B ) = P(A) \cdot P(B) ~ ~ (= P(B) \cdot P(A))\), then they are independent start by checking out that possibility: